4x+2y=-2
3x-4y=-18

Answer by EMStelley(208)

Let's use the addition method here. If we multiply the first equation by 2, we will be able to eliminate y. So we have:
8x + 4y = -4
3x - 4y = -18
If we add these two together we are left with:
11x = -22
So x = -2. If we substitute this back into the original 1st equation, we can find y:
4(-2)+ 2y = -2
-8 + 2y = -2
2y = 6
y=3
So the solution is (-2,3).